Two rectangles have the same area of 480 cm^2.the diffrence between their lengths is 10cm and the diffrence between their widths is 4cm . calculate the dimension of the rectangles
Answer by stanbon(75887)

Two rectangles have the same area of 480 cm^2.the diffrence between their lengths is 10cm and the difference between their widths is 4cm . calculate the dimensions of the rectangles
----
Dimensions of 1st rectangle: L,W
Dimensions of 2nd rectangle: L+10; W+4
===
LW = 480
(L-10)(W+4) = 480
---
LW = LW+4L - 10W - 40
4L = 10W +40
2L = 5W + 20
L = (5/2)W + 10
---
But LW = 480
So, W = 480/L
---
L = (5/2)(480/L)+10
L = 1200/L + 10
L^2 - 10L - 1200 = 0
(L-40)(L+30) = 0
Positive solution:
L = 40 cm (length)
Then 480/L = W = 120 cm (width)
=====================================
Cheers,
Stan H.